Phil is 28 years older than Tom. Six years ago the sum of their age was 36. How old are they now?
Answer by ikleyn(52794)

.
From the condition, you have these two equations in 2 unknowns


P = T + 28              (1)
(P-6) + (T-6) = 36.     (2)


Simplify eq(2) :


P + T = 36 + 6 + 6,   or

P + T = 48              (2')


From eq(1), substitute expression for P into eq(2'), replacing (excluding) P. You will get


(T + 28) + T = 48.


Simplify and solve for T:


2T = 48 - 28 = 20  ====>  T = 20/2 = 10.


Answer.  Tom is 10 years old.  Phil is 10 + 28 = 38 years old.


Check on your own that all conditions of the problem are satisfied.

Solved.

----------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.